OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, -3) and has a slope of 1/2?

OpenStudy (p0sitr0n):

y=mx+b m=1/2 y=x/2+b replace (-2,-3) -3=-2/2+b -3=-1+b -2=b y=x/2-2

OpenStudy (anonymous):

Thanks!! Can you help me with some more?

OpenStudy (anonymous):

@P0sitr0n